# Interpreting the Hi $21$ cm cosmology maps through Largest Cluster Statistics. Part II. Impact of the realistic foreground and instrumental noise on synthetic SKA1-Low observations

### v2

@@ 1 Introduction
It is generally accepted that conclusions from these image-based methods depend on detecting a large number of Hii regions across different stages and sizes. However, the study by Bag et al. [29, 30] demonstrates that identifying only the largest ionized region (LIR) is sufficient to infer the percolation process. To reach this conclusion, they introduced a novel statistic named Largest Cluster Statistics (LCS), along with a shape-finding algorithm. The LCS analysis of the redshifted 21 cm maps has been demonstrated to be an efficient and robust method for tracking the time evolution of the LIRs during the EoR. The LCS can constrain the reionization model and history by quantifying the morphology of neutral hydrogen distribution during the different stages of the EoR. Specifically, the percolation transition of ionized regions, quantified and constrained via LCS, provides a crucial insight about the underlying reionization model [32]. Our previous work, Dasgupta et al. [33], demonstrates how the convolution of the synthesized beam of the radio interferometric arrays, e.g. SKA1-Low with the target signal, can affect our conclusion about the percolation of Hii regions during reionization. They showed that the apparent percolation transition of ionized regions shifted towards the later stage of reionization, depending upon the array synthesized beam of SKA1-Low and thresholding formalism used in noisy data.
@@
Motivated by this, we present an optimal thresholding strategy to minimize the bias in the recovered percolation transition. We further assess the robustness of our LCS analysis of the 21 cm maps under various foreground contamination scenarios. For this purpose, we consider two sources of foreground contamination: a) extra-galactic point sources and b) diffuse synchrotron and free-free emission. For all of these sky models, we consider the antenna-based gain calibration errors to estimate the maximum tolerance level to recover the reionization history in a robust and relatively unbiased manner using LCS. In addition, we have also investigated the impact of instrumental noise for SKA1-Low on this analysis via the synthetic observations simulated by the 21cmE2E pipeline [^2] https://gitlab.com/samit-pal/21cme2e. This analysis is done by diagnostic tool SURFGEN2 [29, 30] to estimate the LCS and gain insights into the percolation of Hii regions during the EoR.
@@
This paper is organized as follows: In Section 2, we discuss the simulation of the radio sky. Section 3 describes our end-to-end simulation and the methodology used to incorporate antenna-based gain calibration errors and instrumental noise on synthetic SKA1-Low observations. The analysis formalism is presented in Section 4, followed by the results in Section 5. Finally, we summarize and discuss our findings in Section 6. We used best-fitted cosmological parameters from the WMAP five-year data release that have been used throughout the paper, which details as follows: $h=0.7,\Omega_{m}=0.27,\Omega_{\Lambda}=0.73,\Omega_{b}h^{2}=0.0226$ [41].
@@ 2 Simulations of the radio sky
We investigate the robustness of LCS analysis to study the evolution of the largest ionized region during different stages of reionization. This analysis uses synthetic SKA1-Low observations generated from a realistic simulation based on 21cmE2E-pipeline. This section describes the simulation of sky models. The sky models consist of the Hi signal and astrophysical foreground within redshift range $7.2<z<8.8$, corresponding to frequency $\sim 144-173$ MHz.
@@ 2.1 Seminumerical simulation of reionization
In this section, we provide a brief review of the simulation of the Hi fields at different stages of the EoR. For detailed information, readers can refer to Section 2 of Dasgupta et al. [33]. To construct the brightness temperature maps of Hi 21 cm signal, we used the ReionYuga simulation [42, 43, 44]. This simulation employs a semi-numerical approach based on excursion set formalism. The ReionYuga utilizes an N-body simulation to create the distribution of dark matter at a given redshift. Next, a Friends-of-Friends (FoF) halo finding algorithm was used to detect the occurrence of the collapse of dark-matter halos inside this distribution of matter. The first light sources, which emit reionizing photons, are formed halos. The ionization fields created via excursion set formalism are thereafter transformed into the field of $21$ cm brightness temperature. For our analysis, we used the existing simulated Hi $21$ cm maps from Dasgupta et al.[33]. These maps are coeval boxes, where each box measures $143.36$ cMpc on each side and is distributed over a mesh consisting of a $256^{3}$ grid volume. A detailed study of the evolution of the LCS along a lightcone to understand how the lightcone effect biases the percolation curve and affects the distinguishability of source models is presented in Potluri et al. (in prep.).
@@ 2.2 Astrophysical foregrounds
One of the major contaminants of CD/EoR experiments is the astrophysical foreground. Their brightness is 454-5 orders of magnitude higher than the faint $21$ cm signal. The primary components in the foreground include diffuse galactic synchrotron radiation, galactic and extra-galactic free-free radiation and extra-galactic point sources. The foreground emission is expected to be spectrally smooth. However, calibration and instrumental effects introduce additional unsmooth structures. We considered two sources of foreground contamination, diffuse emission and extra-galactic point sources, to test these impacts on the residual maps. The foreground contributions are detailed below.
@@ 2.2.1 Diffuse emission
The diffuse galactic emissions are dominated by the synchrotron and free-free emissions. Being large-scale structures, these are sensitive to shorter baselines. The diffuse synchrotron and free-free emission foregrounds were simulated using models outlined in [45, 46]. The Planck Sky Model (PSM) at $217$ GHz are used to simulate the diffuse maps. At each pixel $p$, the brightness of these emissions is quantified using brightness temperature $T_{s}$. These are modelled as power laws:
@@
where $\beta_{s}$ is the spectral index. The all-sky amplitudes for synchrotron and free-free emissions simulations are publically available on the Planck Legacy Archive [^3] https://pla.esac.esa.int/#maps. Using the healpix routines, we degrade and smooth these maps to our desired nside [^4] The resolution of the map is defined by the nside parameter and resolution. The amplitude of diffuse emission is determined from the PSM at $217$ GHz with a resolution of nside= 2048. To determine the spectral index map of synchrotron emission, we used $217$ GHz and $353$ GHz synchrotron maps at nside=2048. Therefore, the spectral index varies in each pixel for synchrotron maps, whereas free-free emission $\beta_{\text{ff}}=-2.13$ [46], is constant in all the pixels. These spectral indices and brightness temperature maps obtained at $217$ GHz are used to extrapolate to the frequencies of interest using the equation 2.1. Here, we simulated the diffused foregrounds at $\alpha=0$ h and $\delta=-30^{\circ}$ field and rotated the centre of the field to our pointing centre.
@@ 3.3 Instrumental noise
To incorporate the thermal noise into the simulation, we downsampled the simulated 21 cm maps by a factor of $16$ along each side. As discussed in our previous study, the synthetic $21$ cm observations require input maps with dimensions of $2^{3n}$ grids. We used a downsampling algorithm available in the python package scikit-image and integrated this method into the 21cmE2E pipeline. This downsampling is essential to reduce the total observation time to achieve the desired SNR level. After downsampling the maps, the grid size of the final maps is $8.96$ cMpc for redshift $z=7.221$, which corresponds to an angular resolution of $3.45$ arcmin and a frequency resolution of $0.54$ MHz. We generated the image cubes using the 21cmE2E pipeline after adding the thermal noise. In order to balance the resolution and the sensitivity of the telescope, two different imaging weighting schemes are used: natural weighting and briggs weighting. The natural weighting scheme provides a relatively higher SNR but a lower resolution compared to the briggs weighting scheme with a robust parameter of $0.4$. The left and middle panel of Figure 4 shows the slices of the brightness temperature field before and after adding the thermal noise with an observation time of $2000$ h at a channel width of $0.54$ MHz, respectively, at neutral fraction $\bar{x}_{\text{HI}}=0.2$. It is seen that more tunnel-like artefacts are visible on the map after introducing thermal noise. These artefacts result from the effect of noise and deconvolution. In the later section 5.3, we discuss how this feature will affect the percolation process. The right panel of Figure 4 shows the recovered Hii regions after applying an optimum thresholding method. In this map, black(white) regions represent recovered ionized (neutral) regions.
@@ 4.1 Binarization of the image cubes
In order to estimate the LCS on the brightness temperature maps, we use a code, SURFGEN2 [63, 64, 29, 30]. SURFGEN2 not only determines the LCS but also helps identify the topological and morphological features of each ionized region using Shapefinders [65].  [^9] Shapefinders are derived from the ratios of the Minkowski functionals. For instance, in three dimensions, the four Minkowski functionals give rise to three shapefinders, with each one representing the extent of a closed surface along one of the three dimensions; see [65] for more details. For an in-depth understanding of the operating principles of the SURFGEN2 code, readers can refer to [64, 30]. The SURFGEN2 code requires thresholding to binarize the neutral and ionized regions of the Hi maps. In an ideal Hi 21 cm brightness temperature field ($\delta T_{b}$), an ionized region is identified by $\delta T_{b}$ equal to zero. For an interferometric observation, we measure only the fluctuations in the $21$ cm signal. Accordingly, the minima of the brightness temperature maps correspond to ionized regions, as expected. However, in real observations, the presence of systematic effects poses a challenge in finding the optimal threshold to distinguish the ionized regions. In our previous work, Dasgupta et al. [33], imposed a gradient-descent method on the histograms of the synthetically observed Hi maps by the SKA1-Low to binarize them. When we add calibration errors and instrumental noise to the Hi maps, finding an optimal threshold becomes a challenging task.
@@ 5.1 Effect of thresholding
This work presents an optimum strategy to reduce the bias in LCS estimation from the simulated 21 cm observation results. Our previous study, Dasgupta et al. [33] used a thresholding method, e.g., gradient descent, to binarize 21 cm maps and estimate the LCS. During the early stages of reionization, the histogram of 21 cm maps is not bimodal. Instead, it exhibits an unimodal distribution with asymmetric tails, as indicated by non-zero skewness [34, 67]. This thresholding method introduces bias for these cases due to its limitations in reaching the local minima in the 21 cm observation maps. The problem becomes more significant when telescope effects, such as the proper implementation of the array synthesized beam of SKA1-Low, are included. Therefore, these limitations fail to binarize the image pixels and lead to biases in the recovered percolation process of reionization.
@@ 5.2 Effect of gain calibration error
Case I: Figure 6 compares the evolution of the obtained LCS with $\bar{x}_{\text{HI}}$ for varying levels of residual calibration errors on the simulated SKA1-Low observational maps. These maps include both unsubtracted point source contamination and the target 21-cm signal. The effect of residual calibration errors on LCS estimation is shown in magenta circle, cyan diamond, and green square curves in figure 6. These results are obtained from the simulated 21 cm observation via 21cmE2E pipeline based on the threshold set by the optimum thresholding method. The blue solid curve represents the original image, a hypothetical scenario without telescope effect and noise, and the corresponding threshold for LCS is set at zero. We observed that at the residual gain calibration, the inaccuracy of $0.02\%$ (illustrated by the red star curve), the obtained LCS from simulated 21 cm observations remains largely unaffected. It closely matches the evolution of the LCS as seen in the obtained LCS from 21 cm observations without any bias due to corruption.
@@ Appendix A Image Binarization
Threshold selection for identifying neutral and ionized pixels is a crucial aspect of EoR 21 cm image analysis. This work introduces a modified method that combines a global thresholding strategy with an unsharp masking technique, to identify ionized regions within 21cm observation maps. As illustrated in Figure [Figure Number], we compare our method with the binarization technique described by Giri et al. [66]. We find that, whereas their method tends to over-segment ionized regions, particully in areas with low-density fluctuations. Our method demonstrates superior robustness across diverse density environments and more accurately preserves the morphological structure of ionized regions. The black (white) regions in binary maps represent the ionized and neutral regions.

### v3

@@ 1 Introduction
It is generally accepted that conclusions from these image-based methods depend on detecting a large number of Hii regions across different stages and sizes. However, the study by Bag et al. [29, 30] demonstrates that identifying only the largest ionized region (LIR) is sufficient to infer the percolation process. To reach this conclusion, they introduced a novel statistic named Largest Cluster Statistics (LCS), along with a shape-finding algorithm. The LCS analysis of the redshifted 21 cm maps has been demonstrated to be an efficient and robust method for tracking the time evolution of the LIRs during the EoR. The LCS can constrain the reionization model and history by quantifying the morphology of neutral hydrogen distribution during the different stages of the EoR. Specifically, the percolation transition of ionized regions, quantified and constrained via LCS, provides a crucial insight about the underlying reionization model [32]. Our previous work, Dasgupta et al. [33], demonstrated how the convolution of the synthesized beam of the radio interferometric arrays, e.g. SKA1-Low with the target signal, can affect our conclusion about the percolation of Hii regions during reionization. They showed that the apparent percolation transition of ionized regions shifted towards the later stage of reionization, depending upon the array synthesized beam of SKA1-Low and thresholding formalism used in noisy data.
@@
Motivated by this, we present an optimal thresholding strategy to minimize the bias in the recovered percolation transition. We further assess the robustness of our LCS analysis of the 21 cm maps under various foreground contamination scenarios. For this purpose, we consider two sources of foreground contamination: a) extra-galactic point sources and b) diffuse synchrotron and free-free emission. For all of these sky models, we consider the antenna-based gain calibration errors to estimate the maximum tolerance level needed to recover the reionization history in a robust and relatively unbiased manner using LCS. In addition, we have also investigated the impact of instrumental noise for SKA1-Low on this analysis via the synthetic observations simulated by the 21cmE2E pipeline [^2] https://gitlab.com/samit-pal/21cme2e. This analysis is done by diagnostic tool SURFGEN2 [29, 30] to estimate the LCS and gain insights into the percolation of Hii regions during the EoR.
@@
This paper is organized as follows: In Section 2, we discuss the simulation of the radio sky. Section 3 describes our end-to-end simulation and the methodology used to incorporate antenna-based gain calibration errors and instrumental noise on synthetic SKA1-Low observations. The analysis formalism is presented in Section 4, followed by the results in Section 5. Finally, we summarize and discuss our findings in Section 6. We used the best-fitted cosmological parameters from the WMAP five-year data release that have been used throughout the paper, which details as follows: $h=0.7,\Omega_{m}=0.27,\Omega_{\Lambda}=0.73,\Omega_{b}h^{2}=0.0226$ [41].
@@ 2 Simulations of the radio sky
We investigate the robustness of LCS analysis to study the evolution of the largest ionized region during different stages of reionization. This analysis uses synthetic SKA1-Low observations generated from a realistic simulation based on 21cmE2E-pipeline. This section describes the simulation of sky models. The sky models consist of the Hi signal and the astrophysical foreground within redshift range $7.2<z<8.8$, corresponding to frequencies $\sim 144-173$ MHz.
@@ 2.1 Seminumerical simulation of reionization
In this section, we provide a brief review of the simulation of the Hi fields at different stages of the EoR. For detailed information, readers can refer to Section 2 of Dasgupta et al. [33]. To construct the brightness temperature maps of Hi 21 cm signal, we used the ReionYuga simulation [42, 43, 44]. This simulation employs a semi-numerical approach based on the excursion set formalism. The ReionYuga utilizes an N-body simulation to create the distribution of dark matter at a given redshift. Next, a Friends-of-Friends (FoF) halo finding algorithm was used to detect the occurrence of the collapse of dark-matter halos inside this distribution of matter. The first light sources, which emit reionizing photons, are formed halos. The ionization fields created via excursion set formalism are thereafter transformed into the field of $21$ cm brightness temperature. For our analysis, we used the existing simulated Hi $21$ cm maps from Dasgupta et al.[33]. These maps are coeval boxes, where each box measures $143.36$ cMpc on each side and is distributed over a mesh consisting of a $256^{3}$ grid volume. A detailed study of the evolution of the LCS along a lightcone to understand how the lightcone effect biases the percolation curve and affects the distinguishability of source models is presented in Potluri et al. (in prep.).
@@ 2.2 Astrophysical foregrounds
One of the major contaminants of CD/EoR experiments is the astrophysical foreground. Its brightness is 454-5 orders of magnitude higher than the faint $21$ cm signal. The primary components in the foreground include diffuse galactic synchrotron radiation, galactic and extra-galactic free-free radiation and extra-galactic point sources. The foreground emission is expected to be spectrally smooth. However, calibration and instrumental effects introduce additional unsmooth structures. We considered two sources of foreground contamination, diffuse emission and extra-galactic point sources, to test these impacts on the residual maps. The foreground contributions are detailed below.
@@ 2.2.1 Diffuse emission
The diffuse galactic emissions is dominated by the synchrotron and free-free emissions. Being large-scale structures, these are sensitive to shorter baselines. The diffuse synchrotron and free-free emission foregrounds were simulated using models outlined in [45, 46]. The Planck Sky Model (PSM) at $217$ GHz are used to simulate the diffuse maps. At each pixel $p$, the brightness of these emissions is quantified using brightness temperature $T_{s}$. These are modelled as power laws:
@@
where $\beta_{s}$ is the spectral index. The all-sky amplitudes for synchrotron and free-free emissions simulations are publicaly available on the Planck Legacy Archive [^3] https://pla.esac.esa.int/#maps. Using the healpix routines, we degrade and smooth these maps to our desired nside [^4] The resolution of the map is defined by the nside parameter and resolution. The amplitude of diffuse emission is determined from the PSM at $217$ GHz with a resolution of nside= 2048. To determine the spectral index map of synchrotron emission, we used $217$ GHz and $353$ GHz synchrotron maps at nside=2048. Therefore, the spectral index varies in each pixel for synchrotron maps, whereas free-free emission $\beta_{\text{ff}}=-2.13$ [46], is constant in all the pixels. These spectral indices and brightness temperature maps obtained at $217$ GHz are used to extrapolate to the frequencies of interest using the equation 2.1. Here, we simulated the diffused foregrounds at $\alpha=0$ h and $\delta=-30^{\circ}$ field and rotated the centre of the field to our pointing centre.
@@ 3.3 Instrumental noise
To incorporate the thermal noise into the simulation, we downsampled the simulated 21 cm maps by a factor of $16$ along each side. As discussed in our previous study, the synthetic $21$ cm observations require input maps with dimensions of $2^{3n}$ grids. We used a downsampling algorithm available in the python package scikit-image and integrated this method into the 21cmE2E pipeline. This downsampling is essential to reduce the total observation time to achieve the desired SNR level. After downsampling the maps, the grid size of the final maps is $8.96$ cMpc for redshift $z=7.221$, which corresponds to an angular resolution of $3.45$ arcmin and a frequency resolution of $0.54$ MHz. We generated the image cubes using the 21cmE2E pipeline after adding the thermal noise. In order to balance the resolution and the sensitivity of the telescope, two different imaging weighting schemes are used: natural weighting and Briggs weighting. The natural weighting scheme provides a relatively higher SNR but a lower resolution compared to the briggs weighting scheme with a robust parameter of $0.4$. The left and middle panels of Figure 4 shows the slices of the brightness temperature field before and after adding the thermal noise with an observation time of $2000$ h at a channel width of $0.54$ MHz, respectively, at neutral fraction $\bar{x}_{\text{HI}}=0.2$. It is seen that more tunnel-like artefacts are visible on the map after introducing thermal noise. These artefacts result from the effect of noise and deconvolution. In the later section 5.3, we discuss how this feature will affect the percolation process. The right panel of Figure 4 shows the recovered Hii regions after applying an optimum thresholding method. In this map, black(white) regions represent recovered ionized (neutral) regions.
@@ 4.1 Binarization of the image cubes
In order to estimate the LCS on the brightness temperature maps, we use a code, SURFGEN2 [63, 64, 29, 30]. SURFGEN2 not only determines the LCS but also helps identify the topological and morphological features of each ionized region using Shapefinders [65].  [^9] Shapefinders are derived from the ratios of the Minkowski functionals. For instance, in three dimensions, the four Minkowski functionals give rise to three shapefinders, with each one representing the extent of a closed surface along one of the three dimensions; see [65] for more details. For an in-depth understanding of the operating principles of the SURFGEN2 code, readers can refer to [64, 30]. The SURFGEN2 code requires thresholding to binarize the neutral and ionized regions of the Hi maps. In an ideal Hi 21 cm brightness temperature field ($\delta T_{b}$), an ionized region is identified by where $\delta T_{b}$ equal to zero. For an interferometric observation, we measure only the fluctuations in the $21$ cm signal. Accordingly, the minima of the brightness temperature maps correspond to ionized regions, as expected. However, in real observations, the presence of systematic effects poses a challenge in finding the optimal threshold to distinguish the ionized regions. In our previous work, Dasgupta et al. [33], imposed a gradient-descent method on the histograms of the synthetically observed Hi maps from the SKA1-Low to binarize them. When we add calibration errors and instrumental noise to the Hi maps, finding an optimal threshold becomes a challenging task.
@@ 5.1 Effect of thresholding
This work presents an optimum strategy to reduce the bias in LCS estimation from the simulated 21 cm observation results. Our previous study done by Dasgupta et al. [33], used a thresholding method, e.g., gradient descent, to binarize 21 cm maps and estimate the LCS. During the early stages of reionization, the histogram of 21 cm maps is not bimodal. Instead, it exhibits an unimodal distribution with asymmetric tails, as indicated by non-zero skewness [34, 67]. This thresholding method introduces bias for these cases due to its limitations in reaching the local minima in the 21 cm observation maps. The problem becomes more significant when telescope effects, such as the proper implementation of the array synthesized beam of SKA1-Low, are included. Therefore, these limitations fail to binarize the image pixels and lead to biases in the recovered percolation process of reionization.
@@ 5.2 Effect of gain calibration error
Case I: Figure 6 compares the evolution of the obtained LCS with $\bar{x}_{\text{HI}}$ for varying levels of residual calibration errors on the simulated SKA1-Low observational maps. These maps include both unsubtracted point source contamination and the target 21-cm signal. The effect of residual calibration errors on LCS estimation is shown in magenta circle, cyan diamond, and green square curves in figure 6. These results are obtained from the simulated 21 cm observation via 21cmE2E pipeline based on the threshold set by the optimum thresholding method. The blue solid curve represents the original image, a hypothetical scenario without telescope effect and noise, and the corresponding threshold for LCS is set at zero. We observed that with a residual gain calibration inaccuracy of $0.02\%$ (illustrated by the red star curve), the obtained LCS from simulated 21 cm observations remains largely unaffected. It closely matches the evolution of the LCS as seen in the obtained LCS from 21 cm observations without any bias due to corruption.
@@ Appendix A Image Binarization
Threshold selection for identifying neutral and ionized pixels is a crucial aspect of EoR 21 cm image analysis. This work introduces a modified method that combines a global thresholding strategy with an unsharp masking technique, to identify ionized regions within 21cm observation maps. As illustrated in Figure 10, we compare our method with the binarization technique described by Giri et al. [66]. We find that, whereas their method tends to over-segment ionized regions, particully in areas with low-density fluctuations. Our method demonstrates superior robustness across diverse density environments and more accurately preserves the morphological structure of ionized regions. The black (white) regions in binary maps represent the ionized and neutral regions.